# Cooperative ISAC Networks: Performance Analysis, Scaling Laws and Optimization

### v1

@@ II-A Cooperative ISAC Network Model
where ${\bf{w}}^{c}_{i}$ and ${\bf{w}}^{s}_{i}\in{\cal{C}}^{M_{\mathrm{t}}\times 1}$ are normalized transmit beamforming vectors, i.e., $\|{\bf{w}}^{c}_{i}\|=1$ and $\|{\bf{w}}^{s}_{i}\|=1$. Furthermore, $p^{s}$ and $p^{c}$ represent the transmit power of the S&C signals, and $\mathbf{W}_{i}=\left[p^{c}\mathbf{w}^{c}_{i},p^{s}\mathbf{w}^{s}_{i}\right]\in{\cal{C}}^{M_{\mathrm{t}}\times 2}$ is the transmit precoding matrix of the BS at $\mathbf{d}_{i}$. To avoid the interference between S&C, we adopt zero-forcing (ZF) beamforming for the sake of making the analysis tractable. Then, the beamforming vector of the serving BS $i$ is given by
@@ II-B Cooperative Sensing Model
The location of the typical target is denoted by ${\psi}_{t}=[x_{t},y_{t}]^{T}$. Assuming that unbiased measurements can be made, the CRLB is an exemplar benchmark for the theoretical localization accuracy in terms of the mean squared error (MSE) which can be expressed as
${\rm{var}}\{\hat{\psi}_{t}\}={\rm{E}}\{|\hat{\psi}_{t}-{\psi}_{t}|^{2}\}\geq\mathrm{CRLB},\vspace{-1.5mm}$ (3)
where $\hat{\psi}_{t}=\left[\hat{x}_{t},\hat{y}_{t}\right]^{T}$ represents the estimated location of the typical target. In general, time synchronization in coherent distributed MIMO radar is an expensive task, but fortunately a synchronization error corresponding to the propagation delay of the cell is adequate for non-coherent processing [34]. Hence, non-coherent MIMO radar is more practical and thus it is assumed in this work. The typical target is collaboratively sensed by $N$ BSs. Let us assume that the transmitted radar signals $\{s_{i}^{s}\}_{i=1}^{N}$ of the BSs in the cooperative sensing cluster are approximately orthogonal for any time delay of interest [35]. [^1] Due to the substantial amount of original echo data, the echo signals received at each BS can be pre-processed (e.g., coherent operations) in a distributed manner, and then send the processing results to the central unit through backhaul links, which can significantly reduce the backhaul overhead. For non-coherent MIMO radar, the base-band equivalent of the impinging signal at receiver $j$ is represented as
@@
where $\beta\geq 2$ is the pathloss exponent between the serving BS and the typical target, $\sigma$ denotes the radar cross section (RCS), $\tau_{i,j}$ is the propagation delay of the link spanning from BS $i$ to the typical target and then to BS $j$. Finally, the term ${\bf{n}}_{l}(t)$ is the additive complex Gaussian noise having zero mean and covariance matrix ${\bm{\Sigma}}=\sigma_{s}^{2}{\bf{I}}_{M_{\rm{r}}}$. In (4), we have ${{\bf{a}}^{H}}(\theta_{n})=[1,\cdots,e^{{j\pi(M_{\mathrm{t}}-1)\cos(\theta_{n})}}]^{T}$, and ${\bf{b}}(\theta_{j})=[1,\cdots,e^{{j\pi(M_{\mathrm{r}}-1)\cos(\theta_{j})}}]$.
By evaluating the range of each monostatic link and bi-static link, the target location can be estimated by maximum likelihood estimation (MLE) [36]. Then, the Fisher information matrix (FIM) of estimating the parameter vector ${\psi}_{t}$ for the non-coherent MIMO radar considered is equal to [34]
@@
In (6) and (7), the angle $\theta_{i}$ is the angle of bearing for the $i$-th transmitter to the target with respect to the horizontal axis. In (5), we have [34]
@@
$\mathrm{CRLB}={\rm{E}}_{\Phi_{b}}\left[\operatorname{tr}\left(\bar{\mathbf{F}}_{N}^{-1}\left(\psi_{t}\right)\right)\right].\vspace{-1.5mm}$ (9)
@@ III-B Performance Gain of Cooperative Sensing
$\frac{{2\sum\nolimits_{i=1}^{N}\!{\sum\nolimits_{j=1}^{N}d_{i}^{-\beta}d_{j}^{-\beta}\left(1+\cos\left({{\theta_{i}}-{\theta_{j}}}\right)\right)}}}{{\!\sum\nolimits_{l=1}^{N}{\!\sum\nolimits_{k=1}^{N}{\!\sum\nolimits_{i\geq k}^{N}{\!\sum\nolimits_{j>(k-i)N+l}^{N}\!{(d_{i}d_{j}d_{l}d_{k})^{-\beta}}}}}{{\!\left({{a_{kl}}{b_{ij}}\!-\!{a_{ij}}{b_{kl}}}\right)}^{2}}}}\!\bigg{]}\!,\vspace{-1.5mm}$
where $d_{i}={\left\|{{{\bf{d}}_{i}}}\right\|}^{-\beta}$. To obtain a more tractable CRLB expression, we resort to a simple yet tight approximation. Then the following conclusion is proved.
@@
$E\left[{{d_{n}}}\right]={{\frac{{\Gamma\left(n+\frac{1}{2}\right)}}{{\lambda_{b}\pi\Gamma(n)}}}}\approx\frac{\sqrt{n}}{\lambda_{b}\pi}.\vspace{-1.5mm}$ (20)
@@
${\rm{CRLB}}\approx\frac{2}{|\zeta|^{2}{{\lambda_{b}^{2}}{\pi^{2}}\sum\nolimits_{l=1}^{N}{\sum\nolimits_{k=1}^{N}{k^{-\frac{\beta}{2}}l^{-\frac{\beta}{2}}}}}}.\vspace{-1.5mm}$ (21)
@@ III-C Cooperation Acceptance Probability for Limited Resource Blocks
According to Lemma 1, the acceptance probability is monotonically increasing with the number of resource blocks $\psi$, and it is monotonically decreasing with both the target-BS density ratio $\mu_{s}$ and the cluster size $N$. We assume that the number of requests received by different BSs is an independent random variable. Then, the expected CRLB under the consideration of the acceptance probability can be expressed as
@@ VII Conclusion
$=$ ${\sum\nolimits_{l=1}^{N}{\sum\nolimits_{k=1}^{N}{\sum\nolimits_{i\geq k}^{N}{\sum\nolimits_{j>(k-i)N+l}^{N}{{{\left({{a_{kl}}{b_{ij}}-{a_{ij}}{b_{kl}}}\right)}^{2}}}}}}}.\vspace{-1.5mm}$
@@
$=$ $\frac{{2{N^{2}}+2N+2{\rm{E}}\left[{\sum\nolimits_{i\neq j}^{N}{\sum\nolimits_{j=1}^{N}{\cos\left({{\theta_{i}}-{\theta_{j}}}\right)}}}\right]}}{{{\rm{E}}\left[{\sum\nolimits_{l=1}^{N}{\sum\nolimits_{k=1}^{N}{\sum\nolimits_{i\geq k}^{N}{\sum\nolimits_{j>(k-i)N+l}^{N}{{{\left({{a_{kl}}{b_{ij}}-{a_{ij}}{b_{kl}}}\right)}^{2}}}}}}}\right]}}$
@@
${\rm{CRLB}}$ $\!=\!{{\rm{E}}_{d}}\!\left[\!{\frac{{2{{\rm{E}}_{\theta}}\left[{\sum\nolimits_{i=1}^{N}{\sum\nolimits_{j=1}^{N}{{D_{ij}}}}}\right]+2{{\rm{E}}_{\theta}}\left[{\sum\nolimits_{i=1}^{N}{\sum\nolimits_{j=1}^{N}{{D_{ij}}\cos\left({{\theta_{i}}-{\theta_{j}}}\right)}}}\right]}}{{{{\rm{E}}_{\theta}}\left[{\sum\nolimits_{l=1}^{N}{\sum\nolimits_{k=1}^{N}{\sum\nolimits_{i\geq k}^{N}{\sum\nolimits_{j>(k-i)N+l}^{N}{{D_{kl}}{D_{ij}}}}}}{{\left({{a_{kl}}{b_{ij}}-{a_{ij}}{b_{kl}}}\right)}^{2}}}\right]}}}\!\right]\!\!=\!{{\rm{E}}_{d}}\!\left[{\frac{{2\sum\nolimits_{i=1}^{N}{\sum\nolimits_{j=1}^{N}{{D_{ij}}}}+2\sum\nolimits_{i=1}^{N}{{D_{ii}}}}}{{2\sum\nolimits_{l=1}^{N}{\sum\nolimits_{k=1}^{N}{\sum\nolimits_{i\geq k}^{N}{\sum\nolimits_{j>(k-i)N+l}^{N}{{D_{kl}}{D_{ij}}}}}}}}}\right]$

### v3

@@ II-A Cooperative ISAC Network Model
where ${\bf{w}}^{c}_{i}$ and ${\bf{w}}^{s}_{i}\in{\mathbb{C}}^{M_{\mathrm{t}}\times 1}$ are normalized transmit beamforming vectors, i.e., $\|{\bf{w}}^{c}_{i}\|=1$ and $\|{\bf{w}}^{s}_{i}\|=1$. Furthermore, $p^{s}$ and $p^{c}$ represent the transmit power of the S&C signals, and $\mathbf{W}_{i}=\left[p^{c}\mathbf{w}^{c}_{i},p^{s}\mathbf{w}^{s}_{i}\right]\in{\mathbb{C}}^{M_{\mathrm{t}}\times 2}$ is the transmit precoding matrix of the BS at $\mathbf{d}_{i}$. To avoid the interference between S&C, we adopt zero-forcing (ZF) beamforming for the sake of making the analysis tractable. Then, the beamforming vector of the serving BS $i$ is given by
@@ II-B Cooperative Sensing Model
The location of the typical target is denoted by ${{\bm{p}}}_{t}=[x_{t},y_{t}]^{T}$. Assuming that unbiased measurements can be made, the CRLB is an exemplar benchmark for the theoretical localization accuracy in terms of the mean squared error (MSE) which can be expressed as
${\rm{var}}\{\hat{{\bm{p}}}_{t}\}={\rm{E}}\{|\hat{{\bm{p}}}_{t}-{{\bm{p}}}_{t}|^{2}\}\geq\mathrm{CRLB},\vspace{-1.5mm}$ (3)
where $\hat{{\bm{p}}}_{t}=\left[\hat{x}_{t},\hat{y}_{t}\right]^{T}$ represents the estimated location of the typical target. In general, time synchronization in coherent distributed MIMO radar is an expensive task, but fortunately a synchronization error corresponding to the propagation delay of the cell is adequate for non-coherent processing [34]. Hence, non-coherent MIMO radar is more practical and thus it is assumed in this work. The typical target is collaboratively sensed by $N$ BSs. Let us assume that the transmitted radar signals $\{s_{i}^{s}\}_{i=1}^{N}$ of the BSs in the cooperative sensing cluster are approximately orthogonal for any time delay of interest [35]. [^1] Due to the substantial amount of original echo data, the echo signals received at each BS can be pre-processed (e.g., coherent operations) in a distributed manner, and then send the processing results to the central unit through backhaul links, which can significantly reduce the backhaul overhead. For non-coherent MIMO radar, the base-band equivalent of the impinging signal at receiver $j$ is represented as
@@
where $\beta\geq 2$ is the pathloss exponent between the serving BS and the typical target, $\sigma$ denotes the radar cross section (RCS), $\tau_{i,j}$ is the propagation delay of the link spanning from BS $i$ to the typical target and then to BS $j$. Finally, the term ${\bf{n}}_{l}(t)$ is the additive complex Gaussian noise having zero mean and covariance matrix ${\bm{\Sigma}}=\sigma_{s}^{2}{\bf{I}}_{M_{\rm{r}}}$. In (4), we have ${{\bf{a}}^{H}}(\theta_{n})=[1,\cdots,e^{{j\pi(M_{\mathrm{t}}-1)\cos(\theta_{n})}}]^{T}$, and ${\bf{b}}(\theta_{j})=[1,\cdots,e^{{j\pi(M_{\mathrm{r}}-1)\cos(\theta_{j})}}]$, where $\theta_{i}$ is the angle of bearing for the $i$-th BS to the target with respect to the horizontal axis.
By evaluating the range of each monostatic link and bi-static link, the target location can be estimated by maximum likelihood estimation (MLE) [36]. Then, the Fisher information matrix (FIM) of estimating the parameter vector ${{\bm{p}}}_{t}$ for the non-coherent MIMO radar considered is equal to [34]
@@
In (5), we have [34]
@@
$\mathrm{CRLB}={\rm{E}}_{\Phi_{b}}\left[\operatorname{tr}\left(\bar{\mathbf{F}}_{N}^{-1}\right)\right].\vspace{-1.5mm}$ (9)
@@ III-B Performance Gain of Cooperative Sensing
$\frac{{2\sum\nolimits_{i=1}^{N}\!{\sum\nolimits_{j=1}^{N}d_{i}^{-\beta}d_{j}^{-\beta}\left(1+\cos\left({{\theta_{i}}-{\theta_{j}}}\right)\right)}}}{{\!\sum\nolimits_{l=1}^{N}{\!\sum\nolimits_{k=1}^{N}{\!\!\sum\nolimits_{i\geq k}^{N}{\!\sum\nolimits_{j>\!{\lceil(k-i)N+l\rceil\!}^{+}}^{N}\!{(d_{i}d_{j}d_{l}d_{k})^{-\beta}}}}}{{\!\left({{a_{kl}}{b_{ij}}\!-\!{a_{ij}}{b_{kl}}}\right)}^{2}}}}\!\bigg{]}\!,\vspace{-1.5mm}$
where $d_{i}={\left\|{{{\bf{d}}_{i}}}\right\|}$ and ${\lceil x\rceil}^{+}=\max(x,1)$. To obtain a more tractable CRLB expression, we resort to a simple yet tight approximation. Then the following conclusion is proved.
@@
$E\left[{{d_{n}}}\right]={{\frac{{\Gamma\left(n+\frac{1}{2}\right)}}{{\sqrt{\lambda_{b}\pi}\Gamma(n)}}}}\approx\sqrt{\frac{{n}}{\lambda_{b}\pi}}.\vspace{-1.5mm}$ (20)
@@
${\rm{CRLB}}\approx\frac{2}{|\zeta|^{2}{{\lambda_{b}^{\beta}}{\pi^{\beta}}\sum\nolimits_{l=1}^{N}{\sum\nolimits_{k=1}^{N}{k^{-\frac{\beta}{2}}l^{-\frac{\beta}{2}}}}}}.\vspace{-1.5mm}$ (21)
@@ III-C Cooperation Acceptance Probability for Limited Resource Blocks
According to Lemma 1, the acceptance probability is monotonically increasing with the number of resource blocks $\psi$, and it is monotonically decreasing with both the target-BS density ratio $\mu_{s}$ and the cluster size $N$. We assume that the number of requests received by different BSs is an independent random variable. Then, with $\beta=2$, the expected CRLB under the consideration of the acceptance probability can be expressed as
@@ VII Conclusion
$=$ ${\sum\nolimits_{l=1}^{N}{\sum\nolimits_{k=1}^{N}{\sum\nolimits_{i\geq k}^{N}{\sum\nolimits_{j>{\lceil(k-i)N+l\rceil}^{+}}^{N}{{{\left({{a_{kl}}{b_{ij}}-{a_{ij}}{b_{kl}}}\right)}^{2}}}}}}}.\vspace{-1.5mm}$
@@
$=$ $\frac{{2{N^{2}}+2N+2{\rm{E}}\left[{\sum\nolimits_{i\neq j}^{N}{\sum\nolimits_{j=1}^{N}{\cos\left({{\theta_{i}}-{\theta_{j}}}\right)}}}\right]}}{{{\rm{E}}\left[{\sum\nolimits_{l=1}^{N}{\sum\nolimits_{k=1}^{N}{\sum\nolimits_{i\geq k}^{N}{\sum\nolimits_{j>{\lceil(k-i)N+l\rceil}^{+}}^{N}{{{\left({{a_{kl}}{b_{ij}}-{a_{ij}}{b_{kl}}}\right)}^{2}}}}}}}\right]}}$
@@
${\rm{CRLB}}$ $\!=\!{{\rm{E}}_{d}}\!\left[\!{\frac{{2{{\rm{E}}_{\theta}}\left[{\sum\nolimits_{i=1}^{N}{\sum\nolimits_{j=1}^{N}{{D_{ij}}}}}\right]+2{{\rm{E}}_{\theta}}\left[{\sum\nolimits_{i=1}^{N}{\sum\nolimits_{j=1}^{N}{{D_{ij}}\cos\left({{\theta_{i}}-{\theta_{j}}}\right)}}}\right]}}{{{{\rm{E}}_{\theta}}\left[{\sum\nolimits_{l=1}^{N}{\sum\nolimits_{k=1}^{N}{\sum\nolimits_{i\geq k}^{N}{\sum\nolimits_{j>\!{\lceil(k-i)N+l\rceil}^{+}\!}^{N}{{D_{kl}}{D_{ij}}}}}}{{\left({{a_{kl}}{b_{ij}}-{a_{ij}}{b_{kl}}}\right)}^{2}}}\right]}}}\!\right]\!\!=\!{{\rm{E}}_{d}}\!\left[{\frac{{2\sum\nolimits_{i=1}^{N}{\sum\nolimits_{j=1}^{N}{{D_{ij}}}}+2\sum\nolimits_{i=1}^{N}{{D_{ii}}}}}{{\!2\sum\nolimits_{l=1}^{N}\!{\sum\nolimits_{k=1}^{N}\!{\sum\nolimits_{i\geq k}^{N}\!{\sum\nolimits_{j>\!{\lceil(k-i)N+l\rceil}^{+}\!}^{N}{{D_{kl}}{D_{ij}}}}}}}}}\right]$